LASER-INDUCED FLUORESCENCE AND DISPERSED FLUORESCENCE SPECTROSCOPY OF THE TROPYL C_{7}H_{7} RADICAL

\underline{ILIAS SIOUTIS}, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210; GYÖRGY TARCZAY, Department of General and Inorganic Chemistry, Eötvös University, P. O. Box 32, H-1518 Budapest 112, Hungary; and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Laser-induced fluorescence (LIF) and dispersed fluorescence (DF) spectra of the cycloheptatrienyl (tropyl) radical (C_{7}H_{7}) have been observed under supersonic jet cooling conditions. The DF emission was collected by pumping different vibronic bands of the Ã^{2}E′′_{3} ← X̃^{2}E′′_{2} LIF excitation spectrum. Analysis of the DF spectra yields valuable information about the vibrational levels of the ground X̃^{2}E′′_{2} electronic state. In addition to the experimental work, ab initio calculations have been performed providing ground and excited state Jahn-Teller parameters; a comprehensive analysis reveals Jahn-Teller activity in 4 e′_{3} modes for the X̃^{2}E′′_{2} state and 3 e′_{1} modes for the Ã^{2}E′′_{3} and provides values for their vibrational frequencies as well as linear Jahn-Teller constants. In addition to the magnitude of the Jahn-Teller distortion, the stabilization energies for both electronic states have been calculated at the CASSCF(7,7) and EOMEA-CCSD levels of theory using different basis sets.